OpenStudy (anonymous):

How do you do this problem? 32x-8/4 (divided by 4) And 24m-6 divided by -3

OpenStudy (anonymous):

the first equation would look like this, right? \[\frac{ 32x-8 }{4 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Same setup for the second equation too

OpenStudy (anonymous):

then you would solve it like this:\[\frac{ 32x-8 }{ 4}*4\] ( the fours cancel each other out) \[32x-8=0\] (add 8 to both sides) \[32x=8\] then divide both sides by 32 to get your answer

OpenStudy (anonymous):

K thank

OpenStudy (anonymous):

you're welcome. You do the other problem the same way.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."